# Demonstration of logical qubits and repeated error correction with better-than-physical error rates

### v2

@@ Abstract
The promise of quantum computers hinges on the ability to scale to large system sizes, e.g., to run quantum computations consisting of more than 100 million operations fault-tolerantly. This in turn requires suppressing errors to levels inversely proportional to the size of the computation. As a step towards this ambitious goal, we present experiments on a trapped-ion QCCD processor where, through the use of fault-tolerant encoding and error correction, we are able to suppress logical error rates to levels below the physical error rates. In particular, we entangled logical qubits encoded in the $[[7,1,3]]$ code with error rates 9.8×9.8$\times$9.8 × to 500×500$\times$500 × lower than at the physical level, and entangled logical qubits encoded in a $[[12,2,4]]$ code with error rates 4.7×4.7$\times$4.7 × to 800×800$\times$800 × lower than at the physical level, depending on the judicious use of post-selection. Moreover, we demonstrate repeated error correction with the $[[12,2,4]]$ code, with logical error rates below physical circuit baselines corresponding to repeated $\rm CNOT$s, and show evidence that the error rate per error correction cycle, which consists of over 100 physical $\rm CNOT$s, approaches the error rate of two physical $\rm CNOT$s. These results signify an important transition from noisy intermediate scale quantum computing to reliable quantum computing, and demonstrate advanced capabilities toward large-scale fault-tolerant quantum computing.
@@
Quantum computers have the potential to solve important classically-intractable problems, however doing so requires improving error rates well beyond those of the underlying physical hardware. The development of quantum error correction and fault-tolerant quantum computing was a major theoretical breakthrough that paved the way for the implementation of reliable quantum computers [1, 2, 3, 4, 5, 6, 7]. Without quantum fault-tolerance, there is little to no indication that quantum computers can solve important practical problems that are outside the reach of modern day supercomputers and machine learning [8, 9]. The experimental challenges remain significant, as fault tolerance requires that physical error rates be sufficiently low before the overhead of error correction leads to an improvement over physical, non-fault-tolerant operations [3, 5, 6, 7, 10], but steady progress has been made across several different platforms. In fact, several experiments have shown indications of physical error rates approaching this important so-called threshold [11, 12, 13, 14], while others have demonstrated operations on multiple logical qubits [15, 16, 17]. However, to the best of our knowledge none of these experiments have demonstrated logical error rates better than the physical error rates— a notable exception being a demonstration of Bell correlations that are stronger than physical correlations [18].
@@ I Methodology
Our approach largely builds on Gottesman’s proposal in Ref. 22. Namely, we benchmark complete quantum circuits  [^1] Complete quantum circuits are circuits that prepare qubits in a fixed state, perform a sequence of gates, and measure one or more qubits to yield classical output bits., and contrast the error rates of the classical outputs of these circuits—a comparison between the outputs of the unencoded physical circuit to that of the corresponding fault-tolerantly encoded circuit on the same hardware. We deviate from Gottesman’s proposal in two ways – we consider a different metric for the comparison of quantum circuits, and we allow for slightly more general state preparations.
@@ II Hardware platform
All reported demonstrations here were performed on Quantinuum’s H2 trapped-ion processor. The device was recently reported on in detail in Ref. 21, and we give a brief overview here.
@@
The H2 processor is commercially available through both Quantinuum [28] and Azure Quantum [29]. All experiments in Sections III.2 and IV were submitted through the Azure Quantum software stack using the Quantum Intermediate Representation (QIR) [30]. For the experiments described in Section IV, additional compiler customization were applied, as described in that section in more detail.
@@ III.1.1 Circuits
The circuit components used to generate a high-fidelity Bell state were previously demonstrated in Refs. 12, 18. The logical program to prepare a logical Bell resource state using the Steane code is in Fig. 1. The preparation includes encoding circuits to initialize two logical qubits to |$\ket{0}$, transversal single and two-qubit Clifford gates, flagged syndrome extraction, and destructive logical measurements. Each logical qubit has seven data qubits and three ancilla qubits, leading to the experiments having a total of 20 physical qubits.
@@
After running syndrome extraction, logical single-qubit transversal gates are applied to measure in appropriate logical Pauli bases, and then we destructively measure the data qubits. We measure both qubits in the logical $X$, $Y$, and $Z$ bases, which can be performed transversally in the Steane code. The destructive measurements of data qubits not only allows one to determine logical outcome but can also be used to determine syndromes. At the end of each destructive logical measurement, these syndromes are used to generate a correction to the logical outcome using a lookup table decoder. These corrected outcomes are determined by running the decoder during the hybrid quantum/classical program on the device and not determined afterwards. The lookup table is relatively simple and the same one used in Ref. 12. Syndromes are decoded independently for each logical measurements. Further improvements could be made by incorporating experimental bias noise in the construction of the decoder, as well as decoding over all syndromes generated by both logical qubits [38, 39, 40, 17].
The $[[12,2,4]]$ code also studied in this work does not admit for a transversal fault-tolerant $Y$ eigenbasis measurement, so to facilitate comparison we focus on the error rate for Bell experiments where we measure only $X$ and $Z$ parities, and denote it $E_{xz}$. Simply put, $E_{xz}$ is defined as the number of “incorrect” measurement results divided by the total number of trials. Only measuring in the $X$ and $Z$ basis is related to measuring the stabilizers of the target state. These measurements can provide a bound on the fidelity of the operation, but since we can also measure in the $Y$ basis when using the Steane code, we also report estimates of the process fidelity in Appendix C. The derived process fidelities give a more conservative estimate of the overall performance of the Bell preparation protocol; however, the process fidelities paint a similar picture to the estimates given by $E_{xz}$, where the encoded circuits are statistically significantly out performing the results of the unencoded circuit analog.
Besides evaluating $E_{xz}$ in the presence of quantum error correction, we also re-analyze the corrected outcomes calculated by the device by excluding any results for which non-trivial syndromes were recorded during the destructive measurement phase and the evaluate the Steane code as a quantum error detecting (QEC) code. This mode of operation does not satisfy the requirement that all single faults must be corrected, but allows us to evaluate the probability of the environment applying pure logical errors (thus serving as an upper bound on QEC performance) as well as evaluate the rejection rate the cost of using the Steane code in a QEC manner and whether that cost is worth paying to further suppress noise.
@@ III.1.2 Experimental results
The experimental results for the both the Steane code and physical level Bell state preparation are summarized in Table 1 and Fig. 2 (see Appendix A for details of the statistical analysis and Appendix C for additional data).
We ran a total of $411,600$ unencoded experiments (four Bell state preparation and measurement circuits per program). When restricted to $X$ and $Z$ basis experiments, $274,400$ experiments were ran at the unencoded level, respectively. For both the unencoded and encoded set of experiments, we ran an equal number of sub-experiments measuring $X$, $Y$, and $Z$ correlations. Of the $411,600$ physical experiments, $1,897$ measured the incorrect parity. In particular, the wrong parity was measured $572$ out of $137,200$ times for $X$ parity, $530$ out of $137,200$ times for $Y$ parity, and $795$ out of $137,200$ times for $Z$ parity. This results in an error rate $E_{xz}=0.50\%_{-0.03\%}^{+0.03\%}$. Looking at the outcomes for separate bases (see Table 4), it is apparent that the $Z$ measurements experience an increased error rate of approximately $0.6\%$ compared to about $0.4\%$ for the $X$ and $Y$. This difference may arise due to biased noised in the two-qubit gates.
@@
Table 1: Summary of experimental results for the preparation of Bell resource states using the $[[7,1,3]]$ Steane code and the corresponding circuit composed of physical qubits instead of logical qubits. Note this table only includes data for the $X$ and $Z$ basis measurements. Additional data including calculated process fidelities are given in Appendix C. We define “gain” to be the error rate of the physical circuits divided by the error rate of the corresponding logical circuit, while “corrections” refers to the number of pre-accepted experiments where a non-trivial syndrome was observed.
@@ III.2 Carbon code
We also entangled two pairs of logical qubits using a $[[12,2,4]]$ CSS code, which we nickname the Carbon code  [^2] Details of the code and construction of logical operations will be made available in Ref. 49., and perform Bell correlation experiments with this entangled state. As a CSS code, Carbon has transversal $\rm CNOT$ and $\rm H$ gates, which immediately enable the preparation of a maximally entangled logical Bell state. However, the transversal $\rm CNOT$ gates create entanglement between qubit blocks, instead of entanglement within the block. For that reason, the Bell correlation experiment now involves two logical Bell states, as illustrated in Fig. 3.
@@ III.2.1 Circuits
The Carbon code circuit for logical Bell state preparation requires 30 physical qubits (24 data qubits for the two blocks and 6 ancillas). While it is possible to fault-tolerantly prepare each block in tensor products of $X$ or $Z$ eigenstates and then prepare Bell states by applying the transversal $\rm CNOT$ between them, it is more favorable to distribute the verification measurements throughout the Bell state preparation circuit (not shown). These verification circuits are tailored to the Bell state preparation, and can avoid measuring all stabilizer generators, focusing instead of simply detecting the propagation of failures into high weight Pauli errors. This simplification improves the fidelity of accepted state preparations in the experiments at the cost of mildly increasing the rejection rate.
@@
The $X$ and $Z$ parities for each Bell pair can be obtained by measuring each block transversally, measuring both qubits within each block in the same basis  [^3] The measurement of $Y$ parities requires more complex circuitry (effectively applying the $S$ gate to change bases), so we leave these more complex experiments for future work.. As in Section III.1, no error correction is performed between state preparation and measurement, but syndrome information from the transversal measurements allows us to detect errors nonetheless. Because the Carbon code has distance 4, while we can correct all weight 1 errors in each block, we cannot correct all weight 2 errors. However, we can detect all errors of weight up to 2, correct some of them with high probability, and reject the ones we cannot correct with high probability. Thus, the Carbon code experiments can have two modes of operation: an error correction mode, and an error correction and rejection mode. Importantly, error correction is performed in both modes of operation.
@@
In the second mode of operation, we perform pre-selection of state preparation as before, but we also post-select on syndrome outcomes. Post-selection is based on the expected quality of decoding decisions. We modify the decoder table construction so that if two Pauli errors of a particular weight have the same syndrome and their product is not in the stabilizer group, the syndrome is added to the rejection set (so those outcomes are post-rejected), otherwise it is added to the acceptance set. For syndromes in the acceptance set, each syndrome is associated with the lowest weight Pauli error with that syndrome in the decoding look-up table  [^4] Other rules based on information about the error models bias are certainly possible, but were not considered for these experiments..
The advantage of post-selecting on some syndrome outcomes for even distance codes is that it effectively boosts the distance of the code from $d$ to $d+1$, in the sense that errors of weight $\frac{d}{2}$ do not cause logical errors anymore (as those are rejected in post-selection) [44]. Although post-selection is not scalable in a strict sense, with low rejection rates one can run relatively deep circuits and achieve significant improvements to the logical error rate [44, 45]. The key observation is that the post-rejection rate is second order in the physical error rate, since weight 1 errors are corrected instead of rejected, and the constant factor is improved by correcting any unambiguous syndrome of an error with weight 2.
@@ III.2.2 Experimental results
For the unencoded experiments, out of the 16,000 runs, 125 yielded the incorrect parity, resulting is a physical error rate of $0.8\%_{-0.1\%}^{+0.1\%}$.
@@
Using the methods described in Appendix A, we estimate the error rate for the unencoded circuit is $0.8\%_{-0.1\%}^{+0.1\%}$, while our estimate for the encoded experiments with pre-selection is $0.17\%_{-0.06\%}^{+0.07\%}$ (a reduction in error rate by a factor of 4.7). Our estimate for the encoded experiments with pre-selection and post-selection is $0.001\%_{-0.001\%}^{+0.015\%}$, corresponding to a reduction in error rate by a factor of 800.
@@ IV Repeated fault-tolerant error correction
It is not sufficient to show shallow logical circuits that outperform their physical counterparts. The promise of quantum computers lies with solving large practical problems that require deep quantum circuits [8, 9]. This, in turn, requires fault-tolerant gadgets that intermix quantum error correction with fault-tolerant logical operations [3, 5, 6, 7]. Composable, repeatable error correction is one of the key milestones in this journey, and only recently experimental demonstrations have become possible [47, 48, 13]. Two notable demonstrations from the last several years deserve special mention. A demonstration for the $[[7,1,3]]$ Steane code achieved $\approx 1.75\%$ logical error rate per error correction cycle for up to 6 rounds of error correction [12], a logical error rate just under one order of magnitude higher than the dominating error rate for the physical operations. Repeated error correction was also demonstrated for the surface code [13], achieving error rates of $\approx 9\%$ to $\approx 15\%$ per error correction cycle for up to 8 rounds of error correction, which is again just under one order of magnitude higher than the dominating error rate for the physical operations.
Here we demonstrate as many as 3 rounds of error correction for the $[[12,2,4]]$ Carbon code, using a combination of error correction and rejection. The Carbon code has a high threshold of approximately $1\%$ and a rate of $\frac{1}{6}$ at distance 4 [49], which makes it a good candidate for experimental demonstrations. We compare the error rates of logical circuits with error correction and physical circuits, and show small circuits with logical error rates lower than the physical error rates by a statistically significant margin. We also give evidence that the error rate accumulated per round of error correction is comparable to the error rate accumulated with two physical $\rm CNOT$ in series.

### v3

@@ Abstract
The promise of quantum computers hinges on the ability to scale to large system sizes, e.g., to run quantum computations consisting of more than 100 million operations fault-tolerantly. This in turn requires suppressing errors to levels inversely proportional to the size of the computation. As a step towards this ambitious goal, we present experiments on a trapped-ion QCCD processor where, through the use of fault-tolerant encoding and error correction, we are able to suppress logical error rates to levels below the physical error rates. In particular, we entangled logical qubits encoded in the $[[7,1,3]]$ code with error rates 9.8×9.8$\times$9.8 × to 500×500$\times$500 × lower than at the physical level, and entangled logical qubits encoded in a $[[12,2,4]]$ code based on Knill’s C4/C6 scheme with error rates 4.7×4.7$\times$4.7 × to 800×800$\times$800 × lower than at the physical level, depending on the judicious use of post-selection. Moreover, we demonstrate repeated error correction with the $[[12,2,4]]$ code, with logical error rates below physical circuit baselines corresponding to repeated $\rm CNOT$s, and show evidence that the error rate per error correction cycle, which consists of over 100 physical $\rm CNOT$s, approaches the error rate of two physical $\rm CNOT$s. These results signify a transition from noisy intermediate scale quantum computing to reliable quantum computing, and demonstrate advanced capabilities toward large-scale fault-tolerant quantum computing.
@@
Quantum computers have the potential to solve important classically-intractable problems, however doing so requires improving error rates well beyond those of the underlying physical hardware. The development of quantum error correction and fault-tolerant quantum computing was a major theoretical breakthrough that paved the way for the implementation of reliable quantum computers [1, 2, 3, 4, 5, 6, 7]. Without quantum fault-tolerance, there is little to no indication that quantum computers can solve important practical problems that are outside the reach of modern day supercomputers and machine learning [8, 9]. The experimental challenges remain significant, as fault tolerance requires that physical error rates be sufficiently low before the overhead of error correction leads to an improvement over physical, non-fault-tolerant operations [3, 5, 6, 7, 10]. Steady progress has been made across several different platforms. Several experiments have shown indications of physical error rates approaching this important so-called threshold [11, 12, 13, 14, 15], while others have demonstrated operations on multiple logical qubits [16, 17, 18]. However, to the best of our knowledge none of these experiments have convincingly demonstrated error correction with logical error rates better than the physical error rates.  [^1] A notable exception being a demonstration of Bell correlations that are stronger than physical correlations [15]. Since the original writing, other notable demonstrations have occured, including definitive sub-threshold performance [55], and additional better-than-physical performance [56].
@@ I Methodology
Our approach largely builds on Gottesman’s proposal in Ref. 23. Namely, we benchmark complete quantum circuits  [^2] Complete quantum circuits are circuits that prepare qubits in a fixed state, perform a sequence of gates, and measure one or more qubits to yield classical output bits., and contrast the error rates of the classical outputs of these circuits—a comparison between the outputs of the unencoded physical circuit to that of the corresponding fault-tolerantly encoded circuit on the same hardware. We deviate from Gottesman’s proposal in two ways – we consider a different metric for the comparison of quantum circuits, and we allow for slightly more general state preparations.
@@ II Hardware platform
All reported demonstrations here were performed on Quantinuum’s H2 trapped-ion processor. The device was recently reported on in detail in Ref. 22, and we give a brief overview here.
@@
The H2 processor is commercially available through both Quantinuum [29] and Azure Quantum [30]. All experiments in Sections III.2 and IV were submitted through the Azure Quantum software stack using the Quantum Intermediate Representation (QIR) [31]. For the experiments described in Section IV, additional compiler customizations were applied, as described in that section in more detail.
@@ III.1.1 Circuits
The circuit components used to generate a high-fidelity Bell state were previously demonstrated in Refs. 12, 15. The logical program to prepare a logical Bell resource state using the Steane code is in Fig. 1. The preparation includes encoding circuits to initialize two logical qubits to |$\ket{0}$, transversal single and two-qubit Clifford gates, flagged syndrome extraction, and destructive logical measurements. Each logical qubit has seven data qubits and three ancilla qubits, leading to the experiments having a total of 20 physical qubits.
@@
After running syndrome extraction, logical single-qubit transversal gates are applied to measure in appropriate logical Pauli bases, and then we destructively measure the data qubits. We measure both qubits in the logical $X$, $Y$, and $Z$ bases, which can be performed transversally in the Steane code. The destructive measurements of data qubits not only allows one to determine logical outcomes but can also be used to determine syndromes. At the end of each destructive logical measurement, these syndromes are used to generate a correction to the logical outcome using a lookup table decoder. These corrected outcomes are determined by running the decoder during the hybrid quantum/classical program on the device and not determined afterwards. The lookup table is relatively simple and the same one used in Ref. 12. Syndromes are decoded independently for each logical measurement. Further improvements could be made by incorporating experimental bias noise in the construction of the decoder, as well as decoding over all syndromes generated by both logical qubits [39, 40, 41, 18].
The $[[12,2,4]]$ code also studied in this work does not admit for a transversal fault-tolerant $Y$ eigenbasis measurement, so to facilitate comparison we focus on the error rate for Bell experiments where we measure only $X$ and $Z$ parities, and denote it $E_{xz}$. Simply put, $E_{xz}$ is defined as the number of “incorrect” measurement results divided by the total number of trials. Only measuring in the $X$ and $Z$ basis is related to measuring the stabilizers of the target state. These measurements can provide a bound on the state fidelity, but since we can also measure in the $Y$ basis when using the Steane code, we also report estimates of the state fidelity in Appendix D. The state fidelities give a more conservative estimate of the overall performance of the Bell preparation protocol; however, the state fidelities paint a similar picture to the estimates given by $E_{xz}$, where the encoded circuits are significantly out performing the results of the unencoded circuit analog.
Besides evaluating $E_{xz}$ in the presence of quantum error correction, we also re-analyze the corrected outcomes calculated by the device by excluding any results for which non-trivial syndromes were recorded during the destructive measurement phase and the evaluate the Steane code as a quantum error detecting (QED) code. This mode of operation does not satisfy the requirement that all single faults must be corrected, but allows us to evaluate the probability of the environment applying pure logical errors (thus serving as an upper bound on QEC performance) as well as evaluate the rejection rate cost of using the Steane code in a QED manner and whether that cost is worth paying to further suppress noise.
@@ III.1.2 Experimental results
The experimental results for the both the Steane code and physical level Bell state preparation are summarized in Table 1 and Fig. 2 (see Appendix B for details of the statistical analysis and Appendix D for additional data).
We ran a total of $411,600$ unencoded experiments (four Bell state preparation and measurement circuits per program). When restricted to $X$ and $Z$ basis experiments, $274,400$ experiments were ran at the unencoded level, respectively. For both the unencoded and encoded set of experiments, we ran an equal number of sub-experiments measuring $X$, $Y$, and $Z$ correlations. Of the $411,600$ physical experiments, $1,897$ measured the incorrect parity. In particular, the wrong parity was measured $572$ out of $137,200$ times for $X$ parity, $530$ out of $137,200$ times for $Y$ parity, and $795$ out of $137,200$ times for $Z$ parity. This results in an error rate $E_{xz}=0.50\%_{-0.03\%}^{+0.03\%}$. Looking at the outcomes for separate bases (see Table 5), it is apparent that the $Z$ measurements experience an increased error rate of approximately $0.6\%$ compared to about $0.4\%$ for the $X$ and $Y$. This difference may arise due to biased noise in the two-qubit gates.
@@
Table 1: Summary of experimental results for the preparation of Bell resource states using the $[[7,1,3]]$ Steane code and the corresponding circuit composed of physical qubits instead of logical qubits. Note this table only includes data for the $X$ and $Z$ basis measurements. Additional data including the estimated state fidelity are given in Appendix D. We define “gain” to be the error rate of the physical circuits divided by the error rate of the corresponding logical circuit, while “corrections” refers to the number of pre-accepted experiments where a non-trivial syndrome was observed.
@@ III.2 Carbon code
We also entangled two pairs of logical qubits using a $[[12,2,4]]$ CSS code, which we nickname Carbon (see Appendix A for details), and perform Bell correlation experiments with this entangled state. Carbon is the result of concatenating the $[[4,2,2]]$ code with the $[[6,2,2]]$ code, as proposed originally by Knill [42, 43]. The main modification of Knill’s original proposal is that the circuits here do not obey the concatenated structure, resulting in much more compact state preparation and syndrome extraction circuits. As a CSS code, Carbon has transversal $\rm CNOT$ and $\rm H$ gates, which immediately enable the preparation of a maximally entangled logical Bell state. However, the transversal $\rm CNOT$ gates create entanglement between qubit blocks, instead of entanglement within the block. For that reason, the Bell correlation experiment now involves two logical Bell states, as illustrated in Fig. 3.
@@ III.2.1 Circuits
The Carbon code circuit for logical Bell state preparation requires 30 physical qubits (24 data qubits for the two blocks and 6 ancillas). While it is possible to fault-tolerantly prepare each block in tensor products of $X$ or $Z$ eigenstates and then prepare Bell states by applying the transversal $\rm CNOT$ between them, it is more favorable to distribute the verification measurements throughout the Bell state preparation circuit. See Figs. 9 and 10. These verification circuits are tailored to the Bell state preparation, and can avoid measuring all stabilizer generators, focusing instead of simply detecting the propagation of failures into high weight Pauli errors. This simplification improves the fidelity of accepted state preparations in the experiments at the cost of mildly increasing the rejection rate.
@@
The $X$ and $Z$ parities for each Bell pair can be obtained by measuring each block transversally, measuring both qubits within each block in the same basis  [^3] The measurement of $Y$ parities requires more complex circuitry (effectively applying the $S$ gate to change bases), so we leave these more complex experiments for future work.. As in Section III.1, no error correction is performed between state preparation and measurement, but syndrome information from the transversal measurements allows us to detect errors nonetheless. Because the Carbon code has distance 4, while we can correct all weight 1 errors in each block, we cannot correct all weight 2 errors. However, we can detect all errors of weight up to 2 and reject them, if necessary. Thus, the Carbon code experiments can have two modes of operation: an error correction mode, and an error correction and rejection mode. Error correction is performed in both modes of operation.
@@
In the second mode of operation, we perform pre-selection of state preparation as before, but we also post-select on syndrome outcomes. Post-selection is based on the expected quality of decoding decisions. We modify the decoder table construction so that pairs of circuit faults are no longer accepted. Details are discussed in Appendix A.
The advantage of post-selecting on some syndrome outcomes for even distance codes is that it effectively boosts the distance of the code from $d$ to $d+1$, in the sense that errors of weight $\frac{d}{2}$ do not cause logical errors anymore (as those are rejected in post-selection) [45]. Although post-selection is not scalable in a strict sense, with low rejection rates one can run relatively deep circuits and achieve significant improvements to the logical error rate [45, 46]. The key observation is that the post-rejection rate is second order in the physical error rate, since single faults are corrected instead of rejected.
@@ III.2.2 Experimental results
For the unencoded experiments, out of the 16,000 runs, 125 yielded the incorrect parity, resulting in a physical error rate of $0.8\%_{-0.1\%}^{+0.1\%}$.
@@
Using the methods described in Appendix B, we estimate the error rate for the unencoded circuit is $0.8\%_{-0.1\%}^{+0.1\%}$, while our estimate for the encoded experiments with pre-selection is $0.17\%_{-0.06\%}^{+0.07\%}$ (a reduction in error rate by a factor of 4.7). Our estimate for the encoded experiments with pre-selection and post-selection is $0.001\%_{-0.001\%}^{+0.015\%}$, corresponding to a reduction in error rate by a factor of 800.
@@ IV Repeated fault-tolerant error correction
It is not sufficient to show shallow logical circuits that outperform their physical counterparts. The promise of quantum computers lies with solving large practical problems that require deep quantum circuits [8, 9]. This, in turn, requires fault-tolerant gadgets that intermix quantum error correction with fault-tolerant logical operations [3, 5, 6, 7]. Composable, repeatable error correction is one of the key milestones in this journey, and only recently experimental demonstrations have become possible [48, 49, 13]. Two demonstrations from the last several years deserve special mention. A demonstration for the $[[7,1,3]]$ Steane code achieved $\approx 1.75\%$ logical error rate per error correction cycle for up to 6 rounds of error correction [12], a logical error rate just under one order of magnitude higher than the dominating error rate for the physical operations. Repeated error correction was also demonstrated for the surface code [13], achieving error rates of $\approx 9\%$ to $\approx 15\%$ per error correction cycle for up to 8 rounds of error correction, which is again just under one order of magnitude higher than the dominating error rate for the physical operations.  [^5] We again note demonstrations [55, 56] that appeared since our original writing.
Here we demonstrate as many as 3 rounds of error correction for the $[[12,2,4]]$ Carbon code, using a combination of error correction and detection. The Carbon code has a high threshold of approximately $3\%$ and a rate of $\frac{1}{6}$ at distance 4 (see Appendix A), which makes it a good candidate for experimental demonstrations. We compare the error rates of logical circuits with error correction and physical circuits, and show small circuits with logical error rates lower than the physical error rates by a statistically significant margin. We also give evidence that the error rate accumulated per round of error correction is comparable to the error rate accumulated with two physical $\rm CNOT$ in series.